OpenStudy (anonymous):

1 over 9 times (36-45q)=2 over 3 times (5q+31)

OpenStudy (anonymous):

It's a proportion. a/b = c/d is the same as a(d) =b(c)

OpenStudy (anonymous):

Hmmm...

OpenStudy (anonymous):

so how would that play out in this equation?

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

ah...thanks. got it now.